OpenStudy (anonymous):

the midpoint of uv is (5 -11) the coordinates of one endpoint are u 3,5 find the coordinates of endpoint v

OpenStudy (anonymous):

The midpoint is \[ \frac{u+v}{2}=m \]

OpenStudy (anonymous):

We know \(m=(5,-11)\)

OpenStudy (anonymous):

and \(u=(3,5)\)

OpenStudy (anonymous):

Let \(v=(v_1,v_2)\)

OpenStudy (anonymous):

\[ \frac{3+v_1}{2}=5\\ \frac{5+v_2}{2}=-11 \]

OpenStudy (anonymous):

@spada get it?

OpenStudy (anonymous):

so it that the answer

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

No, you need to solve for \(v_1\) and \(v_2\).

OpenStudy (anonymous):

The answer is \(v=(v_1,v_2)\)

OpenStudy (anonymous):

still confused?

OpenStudy (anonymous):

You don't know how to solve an equation?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio \[ \frac{3+v_1}{2}=5\\ \frac{5+v_2}{2}=-11 \] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

what is v1 and v2

OpenStudy (anonymous):

You don't know how to solve for \(v_1\) in the first equation? Is that what you are telling me?

OpenStudy (anonymous):

No one taught you how to isolate a variable, but they expect you to find midpoints? Is that what I'm hearing?

OpenStudy (anonymous):

yes just what is the answer to the problem?

OpenStudy (anonymous):

No dice.

OpenStudy (anonymous):

?